# Totally odd immersions of complete graphs in graph products

### v1

@@ 3 Cartesian Product
So, we focus our attention on the case when at least one of the two graphs is not bipartite, i.e, has $\mathrm{toi}\geq 3$. An obvious upper bound for $\mathrm{toi}(K_{t}\Square K_{s})$ is the maximum degree of this product graph plus 1. Note that the degree of every vertex in $K_{t}\Square K_{s}$ is $t+s-2$, and hence $\mathrm{toi}(K_{t}\Square K_{s})\leq t+s-1$. In the next theorem, we show that $t+s-1$ is also a lower bound for the toi of the Cartesian product of two graphs $G$ and $H$ whenever one of these has $\mathrm{toi}$ at least 3.
@@
Let $G$ and $H$ be two graphs which have $K_{t}$ and $K_{s}$ as a totally odd strong immersion denoted by $I_{G}\subseteq G$ and $I_{H}\subseteq H$ with the terminals $T_{G}=\{u_{1},u_{2},\dots,u_{t}\}$, and $T_{H}=\{v_{1},v_{2},\dots v_{s}\}$, respectively. Denote by $P_{ii^{\prime}}$ and $Q_{jj^{\prime}}$ the two odd paths from terminals $u_{i}$ to $u_{i^{\prime}}$ and $v_{j}$ to $v_{j^{\prime}}$, in $I_{G}$ and $I_{H}$, respectively.
We claim that $G\Square H$ has $K_{s+t-1}$ as a totally odd immersion with terminals $T_{1}\cup T_{2}$, where
@@
In the following, we define odd paths $M_{(i,j)-(i^{\prime},j^{\prime})}$ in $G\Square H$ joining the terminals $(u_{i},v_{j})$ and $(u_{i^{\prime}},v_{j^{\prime}})$, and show that these paths are pairwise edge-disjoint. We know that $T_{1}$ induces a copy of $H$ in the graph $G\Square H$. So, to connect $(u_{1},v_{i})$ to $(u_{1},v_{i^{\prime}})$, we can use the odd strong immersion path in that copy of $H$. More specifically, for any $1\leq j<j^{\prime}\leq s$, if $Q_{jj^{\prime}}$ is $v_{j}-b_{1}-b_{2}-\dots b_{2\ell}-v_{j^{\prime}}$, then we consider the odd path
@@
Now we need to connect the vertices from $T_{1}$ to the vertices from $T_{2}$. For that, we use the following odd paths
@@
For $2\leq i\leq s$ and $12\leq j\leq t-$ 1 , define the odd path $M_{(1,i)-(j,1)}$ as the path obtained by taking the union of the following three paths,
@@
Since all the odd immersion paths ($P_{ii^{\prime}}$’s and $Q_{jj^{\prime}}$s’) are pairwise edge-disjoint in the respective graphs $G$ and $H$, the new defined paths $M_{(i,j)-(i^{\prime},j^{\prime})}$s’ are also pairwise edge-disjoint. See the figure 1 for an example of the paths in the Cartesian product between two complete graphs ∎
[…]
Figure 1: First paths and last paths of the totally odd immersion of $K_{s+t-1}$ in $K_{s}\Square K_{t}$
@@
Let $G$ and $H$ be graphs containing $K_{3}$ as a totally odd strong immersion with terminals $T_{G}=\{u_{1},u_{2},u_{3}\}$ and $T_{H}=\{v_{1},v_{2},v_{3}\}$, respectively. Let $P_{ii^{\prime}}$ and $Q_{jj^{\prime}}$ be the odd paths (used in the respective totally odd strong immersions) from $u_{i}$ to $u_{i^{\prime}}$ and from $v_{j}$ to $v_{j^{\prime}}$ in $G$ and $H$, respectively, for $1\leq i<i^{\prime}\leq 3$ and $1\leq j<j^{\prime}\leq 3$.
We claim that $G\Square H$ has a $K_{4}$ as a totally odd strong immersion. To prove that we consider the set of terminals $\{(u_{1},v_{1}),(u_{2},v_{1}),(u_{3},v_{1}),(u_{1},v_{2})\}$ and define pairwise edge-disjoint odd paths $M_{(i,j)-(i^{\prime},j^{\prime})}$ in $G\Square H$ joining the terminals $(u_{i},v_{j})$ and $(u_{i^{\prime}},v_{j^{\prime}})$. For any $1\leq i<i^{\prime}\leq 3$, if $P_{ii^{\prime}}$ is $u_{i}-a_{1}-a_{2}-\dots a_{2k}-u_{i^{\prime}}$, then to connect $(u_{i},v_{1})$ to $(u_{i^{\prime}},v_{1})$ we take the path
@@
For the second part of the theorem, we need to show that the Cartesian product of two complete graphs, say with vertex sets $\{x_{1},x_{2},x_{3}\}$ and $\{y_{1},y_{2},y_{3}\}$, does not contain $K_{5}$ as a totally odd strong immersion. Suppose for the contradiction that there exists a $K_{5}$ as a totally odd strong immersion. We claim that either there exists $1\leq i\leq 3$ such that $(x_{i},y_{j})$ is a terminal for all $1\leq j\leq 3$ or there exists $1\leq j\leq 3$ such that $(x_{i},y_{j})$ is a terminal for all $1\leq i\leq 3$.
For the sake of contradiction, we assume that there are no $1\leq i\leq 3$ such that $(x_{i},y_{j})$ is a terminal for all $1\leq j\leq 3$ and no $1\leq j\leq 3$ such that $(x_{i},y_{j})$ is a terminal for all $1\leq i\leq 3$. This is only possible when $\{(x_{k},y_{\ell}),(x_{k},y_{\ell+1}),(x_{k+1},y_{\ell}),(x_{k+1},y_{\ell+1}),(x_{k+2},y_{\ell+2})\}$ is the set of terminals, for some $k,\ell\in\{1,2,3\}$, where the indices are in module $3$ ($2+2,3+1,3+2$ are $1,1$ and $2$, respectively).
@@
$u_{1}$$u_{2}$$u_{3}$$v_{1}$$v_{2}$$v_{3}$$u_{4}$$u_{5}$
Figure 4: A totally odd $K_{4}$ immersion in $C_{5}\Square P_{3}$. Boxed vertices are the terminals considered in the proof. Normal edges are the connections that naturally exist in $C_{5}\Square P_{3}$. Dashed edges represent path $M_{(3,1)-(1,1)}$, while dotted edges represent path $M_{(2,1)-(1,2)}$. Finally, path $M_{(1,2)-(3,1)}$ is depicted in wavy edges.
@@
The maximum degree of a vertex in $K_{t}\Square K_{s}$ is $t+s-2$, and thus, $\mathrm{toi}(K_{t}\Square K_{s})\leq t+s-1$. Therefore, Theorem 3.1 gives $\mathrm{toi}(G\Square H)\geq\mathrm{toi}(K_{t}\Square K_{s})$, when $t>3$ or $s>3$. So we may assume that both $s\leq 3$ and $t\leq 3$. If both $s$ and $t$ are $3$, then by Theorem 3.2, $\mathrm{toi}(G\Square H)\geq\mathrm{toi}(K_{t}\Square K_{s})$. So we may assume that either $s=2$ or $t=2$. Without loss of generality we can assume that $t=2.$ If $s$ is also $2$, then $K_{t}\Square K_{s}$ is a bipartite graph and we have $\mathrm{toi}(G\Square H)\geq\mathrm{toi}(K_{t}\Square K_{s})=2$. The only case remaining is $s=3$ and $t=2$ which follows by Theorem 3.3. ∎
Since $\chi(G\Square H)\leq\max\{\chi(G),\chi(H)\}$, then we can also deduce from Theorems 3.1, 3.2 and 3.3 that there are no minimal counterexamples to Conjecture 1.3 that are the Cartesian product of two graphs.
4 Direct Product
In this section we study the behaviour of the parameter $\mathrm{toi}$ under the operation of direct product. The direct product of a bipartite graph with any other graph is a bipartite graph. Thus, $\mathrm{toi}(G\times H)\leq 2$, if either $G$ or $H$ is bipartite. Thus it is enough to consider $\mathrm{toi}(G\times H)$ when neither $G$ nor $H$ is a bipartite graph.
Theorem 4.1.
$\mathrm{toi}(G\times H)\geq\mathrm{toi}(K_{t}\times K_{s}).$
@@ 4 Direct Product
Let $I_{G}\subseteq G$ and $I_{H}\subseteq H$ be $K_{t}$ and $K_{s}$ totally odd strong immersions in $G$ and $H$, respectively and $T_{G}=\{u_{1},u_{2},\dots,u_{t}\}$ and $T_{H}=\{v_{1},v_{2},\dots,v_{s}\}$ be the respective set of terminals. We denote the vertex sets of $K_{t}$ and $K_{s}$ by $\{a_{1},a_{2},a_{3},\dots,a_{t}\}$ and $\{b_{1},b_{2},b_{3},\dots,b_{s}\}$, respectively. Suppose that $\mathrm{toi}(K_{t}\times K_{s})=r$, i.e, $K_{t}\times K_{s}$ has a $K_{r}$-totally odd immersion, say $I_{K_{t}\times K_{s}}$ with the set of terminals $\{c_{1},c_{2},c_{3},\dots,c_{r}\}$. Note that for every $1\leq k\leq r,c_{k}=(a_{i},b_{j})$, for some $1\leq i\leq t$ and $1\leq j\leq s$.
Now we are going to construct a totally odd strong immersion of $K_{r}$ in the graph $G\times H$. Let us define a function $\phi:K_{t}\times K_{s}\rightarrow I_{G}\times I_{H}$ as $\phi(a_{i},b_{j})=(u_{i},v_{j})$, for all $1\leq i\leq t$ and $1\leq j\leq s$. Now we claim that $G\times H$ has a totally odd strong immersion of $K_{r}$ with set of terminals $\{\phi(a,b)\mid(a,b)\in I_{K_{t}\times K_{s}}\}$.
First, for all $1\leq i<i^{\prime}\leq t$ and $1\leq j<j^{\prime}\leq s$, we connect the $(u_{i},v_{j})$ to $(u_{i^{\prime}},v_{j^{\prime}})$ and $(u_{i},v_{j^{\prime}})$ to $(u_{i^{\prime}},v_{j})$ by odd paths $M_{(i,j)-(i^{\prime},j^{\prime})}$ and $M_{(i,j^{\prime})-(i^{\prime},j)}$, respectively, in $K_{t}\times K_{s}$ Since $i\neq i^{\prime}$ and $j\neq j^{\prime}$, both $(a_{i},b_{j})-(a_{i^{\prime}},b_{j^{\prime}})$ and $(a_{i},b_{j^{\prime}})-(a_{i^{\prime}},b_{j})$ are edges in $K_{t}\times K_{s}$. Let $P_{ii^{\prime}}=u_{i}-x_{1}-x_{2}-\dots-x_{k}-u_{i^{\prime}}$ and $Q_{jj^{\prime}}=v_{j}-y_{1}-y_{2}-\dots-y_{\ell}-v_{j^{\prime}}$ be the odd paths of the immersions $I_{G}$ and $I_{H}$, respectively. In particular, we have that no internal vertex is a terminal vertex and they are pairwise edge-disjoint with the other odd paths of the respective immersions. Now we define the odd path $M_{(i,j)-(i^{\prime},j^{\prime})}$ from $(u_{i},v_{j})$ to $(u_{i^{\prime}},v_{j^{\prime}})$ in $G\times H$ as follows.
If $1<k<\ell$, we define $M_{(i,j)-(i^{\prime},j^{\prime})}$ as
$(u_{i},v_{j})-(x_{1},y_{1})-(x_{2},y_{2})-\dots-(x_{k},y_{k})-(x_{k-1},y_{k+1})-\dots-(x_{k},y_{\ell})-(u_{i^{\prime}},v_{j^{\prime}}),$
where from $(x_{k},y_{k})$ to $(x_{k},y_{\ell})$ the first coordinate alternates between $x_{k}$ and $x_{k-1}$. Note that this is valid path, as $k$ and $\ell$ are of the same parity.
If $1<\ell<k$, then $M_{(i,j)-(i^{\prime},j^{\prime})}$ is
$(u_{i},v_{j})-(x_{1},y_{1})-(x_{2},y_{2})-\dots-(x_{\ell},y_{\ell})-(x_{\ell+1},y_{\ell-1})-\dots-(x_{k},y_{\ell})-(u_{i^{\prime}},v_{j^{\prime}})$
where from $(x_{\ell},y_{\ell})$ to $(x_{k},y_{\ell})$ the second coordinate alternates between $y_{\ell}$ and $y_{\ell-1}$.
We have three more possibilities: i) $k=\ell$, ii) $k=0,\ell>1$ and iii) $k>1,\ell=0$. For each of those possibilities, we define $M_{(i,j)-(i^{\prime},j^{\prime})}$ according to the rules stated in Table 3.
definition of $M_{(i,j)-(i^{\prime},j^{\prime})}$ condition
$(u_{i},v_{j})-(x_{1},y_{1})-(x_{2},y_{2})-\cdots-(x_{k},y_{\ell})-(u_{i^{\prime}},v_{j^{\prime}})$ $k=\ell$
$(u_{i},v_{j})-(u_{i^{\prime}},y_{1})-(u_{i},y_{2})-\dots-(u_{i},y_{\ell})-(u_{i^{\prime}},v_{j^{\prime}})$ $k=0$, $\ell>1$
$(u_{i},v_{j})-(x_{1},v_{j^{\prime}})-(x_{2},v_{j})-\dots-(x_{k},v_{j})-(u_{i^{\prime}},v_{j^{\prime}})$ $\ell=0$, $k>1$
Table 3: Rules for the definition of odd paths joining $(u_{i},v_{j})$ to $(u_{i^{\prime}},v_{j^{\prime}})$ in the case that $1\leq i<i^{\prime}\leq t$ and $1\leq j<j^{\prime}\leq s$.
Now, we have to define the odd path $M_{(i,j^{\prime})-(i^{\prime},j)}$ joining $(u_{i},v_{j^{\prime}})$ to $(u_{i^{\prime}},v_{j})$. If $1<k<l$, we define $M_{(i,j^{\prime})-(i^{\prime},j)}$ as it follows
$(u_{i},v_{j^{\prime}})-(x_{1},y_{\ell})-(x_{2},y_{\ell-1})-\dots-(x_{k},y_{\ell-k+1})-(x_{k-1},y_{\ell-k})--\dots-(x_{k},y_{1})-(u_{i^{\prime}},v_{j}),$
where from $(x_{k},y_{\ell-k+1})$ to $(x_{k},y_{1})$ the first coordinate alternates between $x_{k}$ and $x_{k-1}$.
If $1<\ell<k$, then $M_{(i,j^{\prime})-(i^{\prime},j)}$ is
$(u_{i},v_{j^{\prime}})-(x_{1},y_{\ell})-(x_{2},y_{\ell-1})-\dots-(x_{\ell},y_{1})-(x_{\ell+1},y_{2})-\dots-(x_{k},y_{1})-(u_{i^{\prime}},v_{j})$
where from $(x_{\ell},y_{1})$ to $(x_{k},y_{1})$ the second coordinate alternates between $y_{1}$ and $y_{2}$.
For other possibilities, we define $M_{(i,j^{\prime})-(i^{\prime},j)}$ according to the rules stated in Table 4.
definition of $M_{(i,j^{\prime})-(i^{\prime},j)}$ condition
$(u_{i},v_{j^{\prime}})-(x_{1},y_{1})-(x_{2},y_{2})-\dots-(x_{k},y_{\ell})-(u_{i^{\prime}},u_{j})$. $k=\ell$
$(u_{i},v_{j^{\prime}})-(u_{i^{\prime}},y_{1})-(u_{i},y_{2})-\dots-(u_{i},y_{\ell})-(u_{i^{\prime}},v_{j})$ $k=0$, $\ell>1$
$(u_{i},v_{j^{\prime}})-(x_{1},v_{j})-(x_{2},v_{j^{\prime}})-\dots-(x_{k},v_{j})-(u_{i^{\prime}},v_{j})$ $\ell=0$, $k>1$
Table 4: Rules for the definition of odd paths joining $(u_{i},v_{j^{\prime}})$ to $(u_{i^{\prime}},v_{j})$ in the case that $1\leq i<i^{\prime}\leq t$ and $1\leq j<j^{\prime}\leq s$.
Note that $P_{ii^{\prime}}$s’ and $Q_{jj^{\prime}}$s’ are pairwise edge-disjoint odd paths. Thus, two odd paths obtained from the above method are edge-disjoint (method to argue this is exactly the same we used in the last paragraph of the proof of Theorem 2.1).
Now, we show that $\{\phi(a,b)\mid(a,b)\in I_{K_{t}\times K_{s}}\}$ is the set of terminals of a $K_{r}$-totally odd immersion in $G\Square H$, for which we start giving suitable odd paths connecting two terminals. For any $1\leq i\leq i^{\prime}\leq t$ and $j,j^{\prime}\in\{1,2,\dots s\}$, if $\phi(a_{i},b_{j})(=(u_{i},v_{j}))$ and $\phi(a_{i^{\prime}},b_{j^{\prime}})(=(u_{i^{\prime}},v_{j^{\prime}}))$ are two terminals, then we consider the following odd path connecting $(u_{i},v_{j})$ to $(u_{i^{\prime}},v_{j^{\prime}})$.
  • If $i<i^{\prime}$ and $j\neq j^{\prime}$, then $M_{(i,j)-(i^{\prime},j^{\prime})}$ is the suitable odd path connecting $(u_{i},v_{j})$ to $(u_{i^{\prime}},v_{j^{\prime}})$.
  • If either $i=i^{\prime}$ or $j=j^{\prime}$, then first we consider the odd path
    $(a_{i},b_{j})-(a_{i_{1}},b_{j_{1}})-(a_{i_{2}},b_{j_{2}})-\dots-(a_{i_{q}},b_{j_{q}})-(a_{i^{\prime}},b_{j^{\prime}})$
    joining $(a_{i},b_{j})$ to $(a_{i^{\prime}},b_{j^{\prime}})$ in the totally odd strong immersion of $K_{r}$ in $K_{t}\times K_{s}$. Now we define the odd path of the immersion of $K_{r}$ in $G\times H$ joining $(u_{i},v_{j})$ to $(u_{i^{\prime}},v_{j^{\prime}})$ as the union of the following odd paths
    $M_{(i,j)-({i_{1}},{j_{1}})}-M_{(i_{1},j_{1})-(i_{2},j_{2})}-M_{({i_{2}},{j_{2}})-({i_{3}},{j_{3}})}-\dots-M_{(i_{q-1},j_{q-1})-(i_{q},j_{q})}-M_{(i_{q},j_{q})-(i^{\prime},j^{\prime})}.$
Note, these odd paths are edge-disjoint as $\{M_{(i,j)-(i^{\prime},j^{\prime})}\mid 1\leq i<i^{\prime}\leq t,1\leq j<j^{\prime}\leq s\}$ is a set of pairwise edge-disjoint odd paths in $G\times H$. Therefore, $G\times H$ has $K_{r}$ as a totally odd strong immersion. ∎
If we have $V(K_{t})=\{a_{1},a_{2},a_{3},\dots,a_{t}\}$, and $V(K_{s})=\{b_{1},b_{2},b_{3},\dots,b_{s}\}$, with $t\leq s$, then it is clear that the set $\{(a_{i},b_{i})\in V(K_{t}\times K_{s}):i\in\{1,\dots,t\}\}$ induces a $K_{t}$ subgraph in $K_{t}\times K_{s}$. Since $\chi(G\times H)\leq\min\{\chi(G),\chi(H)\}$ for every pair of graphs $G$ and $H$, and since $\times$ is commutative, this and Theorem 4.1 implies that no minimal counterexample to Conjecture 1.3 is the direct product of two graphs.

### v4

@@ 3 Cartesian Product
So, we focus our attention in the case when at least one of the two graphs is not bipartite, and hence, its $\mathrm{toi}$ value is at least $3$. An obvious upper bound for $\mathrm{toi}(K_{t}\Square K_{s})$ is the maximum degree of this product graph $K_{t}\Square K_{s}$ plus 1. Note that the degree of every vertex in $K_{t}\Square K_{s}$ is t+s2t+s-2, and hence $\mathrm{toi}(K_{t}\Square K_{s})\leq t+s-1$. In the next theorem, we show that t+s1t+s-1 is also a lower bound for $\mathrm{toi}$ of the Cartesian product of two graphs $G$ and $H$ whenever one of these has $\mathrm{toi}$ at least 3.
@@
Let $u_{1},u_{2},\dots u_{t}$ be the terminals of a totally odd strong immersion of $K_{t}$ in $G$ and, for $1\leq i<i^{\prime}\leq t$, let $P_{ii^{\prime}}$ be the corresponding odd path joining $u_{i}$ and $u_{i^{\prime}}$. Analogously, we let $v_{1},v_{2},\dots v_{s}$ be the terminals of a totally odd strong immersion of $K_{s}$ in $H$ and $Q_{jj^{\prime}}$ be the corresponding odd path joining $v_{j}$ and $v_{j^{\prime}}$, where $1\leq j<j^{\prime}\leq s$.
We claim that $G\Square H$ has $K_{s+t-1}$ as a totally odd strong immersion with terminals $T_{1}\cup T_{2}$, where
@@
In the following, we define odd paths $M_{(i,j)-(i^{\prime},j^{\prime})}$ in $G\Square H$ joining the terminals $(u_{i},v_{j})$ and $(u_{i^{\prime}},v_{j^{\prime}})$, and show that these paths are pairwise edge-disjoint. We know that $T_{1}$ induces a copy of $H$ in the graph $G\Square H$. So, to connect $(u_{1},v_{i})$ to $(u_{1},v_{i^{\prime}})$, we can use the totally odd strong immersion path in that copy of $H$. More specifically, for any $1\leq j<j^{\prime}\leq s$, if $Q_{jj^{\prime}}$ is $v_{j}-b_{1}-b_{2}-\dots b_{2\ell}-v_{j^{\prime}}$, then we consider the odd path
@@
Now we need to connect the vertices of $T_{1}$ to the vertices of $T_{2}$. For that, we use the following odd paths
@@
For $2\leq i\leq s$ and $2\leq j\leq t-1$ , define the odd path $M_{(1,i)-(j,1)}$ as the path obtained by taking the union of the following three paths,
@@
Since all the odd immersion paths ($P_{ii^{\prime}}$’s and $Q_{jj^{\prime}}$s’) are pairwise edge-disjoint in the respective graphs $G$ and $H$, the new defined paths $M_{(i,j)-(i^{\prime},j^{\prime})}$s’ are also pairwise edge-disjoint. See the figure 10 for an example of the paths in the Cartesian product between two complete graphs. ∎
Figure 8: When $i=1$ and $j\in\{1,2,\dots s\}$
Figure 9: When $i=t$ and $j\in\{1,2,\dots s\}$
Figure 10: The picture shows the odd paths of the totally odd strong immersion of $K_{t+s-1}$ where the red vertices are its terminals, In particular the odd paths which are described are when $i\in\{2,t\}$ and $j\in\{2,3,\dots,s\}$.
@@
Let $u_{1},u_{2},u_{3}$ be the terminals of a totally odd strong immersion of $K_{3}$ in $G$ and, for $1\leq i<i^{\prime}\leq 3$, let $P_{ii^{\prime}}$ be the corresponding odd path joining $u_{i}$ and $u_{i^{\prime}}$. Analogously, we let $v_{1},v_{2},v_{3}$ be the terminals of a totally odd strong immersion of $K_{3}$ in $H$ and $Q_{jj^{\prime}}$ be the corresponding odd path joining $v_{j}$ and $v_{j^{\prime}}$, where $1\leq j<j^{\prime}\leq 3$.
We claim that $G\Square H$ has a $K_{4}$ as a totally odd strong immersion. To prove this, we consider the set of terminals $\{(u_{1},v_{1}),(u_{2},v_{1}),(u_{3},v_{1}),(u_{1},v_{2})\}$ and define pairwise edge-disjoint odd paths $M_{(i,j)-(i^{\prime},j^{\prime})}$ in $G\Square H$ joining the terminals $(u_{i},v_{j})$ and $(u_{i^{\prime}},v_{j^{\prime}})$. For any $1\leq i<i^{\prime}\leq 3$, if $P_{ii^{\prime}}$ is $u_{i}-a_{1}-a_{2}-\dots a_{2k}-u_{i^{\prime}}$, then to connect $(u_{i},v_{1})$ to $(u_{i^{\prime}},v_{1})$ we take the path
@@
For the second part of the theorem, we need to show that the Cartesian product of two complete graphs of size three, say with vertex sets $\{x_{1},x_{2},x_{3}\}$ and $\{y_{1},y_{2},y_{3}\}$, does not contain $K_{5}$ as a totally odd strong immersion. Suppose for the contradiction that there exists a $K_{5}$ as a totally odd strong immersion. We claim that either there exists $1\leq i\leq 3$ such that $(x_{i},y_{j})$ is a terminal for all $1\leq j\leq 3$ or there exists $1\leq j\leq 3$ such that $(x_{i},y_{j})$ is a terminal for all $1\leq i\leq 3$.
For the sake of contradiction, we assume that there are no $1\leq i\leq 3$ such that $(x_{i},y_{j})$ is a terminal for all $1\leq j\leq 3$ and no $1\leq j\leq 3$ such that $(x_{i},y_{j})$ is a terminal for all $1\leq i\leq 3$. This is only possible when $\{(x_{k},y_{\ell}),(x_{k},y_{\ell+1}),(x_{k+1},y_{\ell}),(x_{k+1},y_{\ell+1}),(x_{k+2},y_{\ell+2})\}$ is the set of terminals, for some $k,\ell\in\{1,2,3\}$, where the indices are in module 33 ($2+2,3+1,3+2$ are $1,1$ and $2$, respectively).
@@
Figure 13: A totally odd strong $K_{4}$ immersion in $C_{5}\Square P_{3}$. Red vertices are the terminals considered in the proof. Black edges are the connections that naturally exist in $C_{5}\Square P_{3}$. Red edges represent path $M_{(3,1)-(1,1)}$, while dark red edges represent path $M_{(2,1)-(1,2)}$. Finally, path $M_{(1,3)-(1,2)}$ is depicted in green.
Corollary 3.4.
[…]
Proof.
First, the degree of every vertex in $K_{t}\Square K_{s}$ is t+s2t+s-2, and thus, $\mathrm{toi}(K_{t}\Square K_{s})\leq t+s-1$. When $s\geq 4\text{ and }t\geq 2$, from Theorem 3.1 we have the lower bound. The other cases come from the Theorems 3.2 and 3.3. ∎
@@
The maximum degree of a vertex in $K_{t}\Square K_{s}$ is t+s2t+s-2, and thus, $\mathrm{toi}(K_{t}\Square K_{s})\leq t+s-1$. Therefore, Theorem 3.1 gives $\mathrm{toi}(G\Square H)\geq\mathrm{toi}(K_{t}\Square K_{s})$, when $t>3$ or $s>3$. So we may assume that both $s\leq 3$ and $t\leq 3$. If both $s$ and $t$ are $3$, then by Theorem 3.2, $\mathrm{toi}(G\Square H)\geq\mathrm{toi}(K_{t}\Square K_{s})$. So we may assume that either $s=2$ or $t=2$. Without loss of generality, we can assume that $t=2.$ If $s$ is also $2$, then $K_{t}\Square K_{s}$ is a bipartite graph and we have $\mathrm{toi}(G\Square H)\geq\mathrm{toi}(K_{t}\Square K_{s})=2$, since bipartite graphs exclude a $K_{3}$ as a totally odd strong immersion. The only case remaining is $s=3$ and $t=2$ which follows by Theorem 3.3. ∎
Corollary 3.6.
The cartesian product of two graphs is not a minimal counterexample to Conjeture 1.3.
@@
Let us assume that there are graphs $G$ and $H$ so that $G\Square H$ is a minimal counterexample to Conjecture 1.3. Then, $G$ and $H$ satisfy that $\mathrm{toi}(G)\geq\chi(G)$ and $\mathrm{toi}(H)\geq\chi(H)$. Suppose that $\mathrm{toi}(G)\geq 3$ and $\mathrm{toi}(H)\geq 3$, since $\chi(G\Square H)\leq\max\{\chi(G),\chi(H)\}\leq\chi(G)+\chi(H)-1\leq\mathrm{toi}(G)+\mathrm{toi}(H)-1$, then from Theorem 3.1, we have $\chi(G\Square H)\leq\mathrm{toi}(G)+\mathrm{toi}(H)-1\leq\mathrm{toi}(G\Square H)$, we obtain a contradiction with the fact that $G\Square H$ is a minimal counterexample to Conjecture 1.3.
Suppose that $\chi(G)\leq\mathrm{toi}(G)=3$ and $\chi(H)\leq\mathrm{toi}(H)=3$, then $\chi(G\Square H)\leq 3$, but from the Theorem 3.2 $\mathrm{toi}(G\Square H)\geq 4\geq\chi(G\Square H)$, contradiction. Now, without loss of generality, suppose that $\chi(G)\leq\mathrm{toi}(G)=3$ and $\chi(H)\leq\mathrm{toi}(H)=2$, then clearly $\chi(G\Square H)\leq 3$ and from the Theorem 3.3 we get a contradiction. Finally, the case where $\chi(G)\leq\mathrm{toi}(G)=2$ and $\chi(H)\leq\mathrm{toi}(H)=2$ is trivial as mentioned at the beginning of Section 3. ∎